OpenStudy (anonymous):

Find the price per bottle of water if 12 bottles of water cost $4.49 ?

OpenStudy (anonymous):

It will be \[\frac{ 12 }{ 4.49}=\frac{ 1 }{ x }\] now cross multiply them \[12x=4.49*1\]\[x=\frac{ 4.49 }{ 12 }=0.37\] Therefore per bottle the cost would be 37cents.